OpenStudy (anonymous):

if the sides of a triangle are 5,12 and n write an equality that expresses the value n may have

OpenStudy (anonymous):

Is n the hypotenuse or not?

OpenStudy (anonymous):

If it is, use c^2=a^2+b^2. If not, use a^2=c^2-b^2 or b^2=c^2-a^2

OpenStudy (anonymous):

sqaure root n^2 and the other side to solve for n

OpenStudy (anonymous):

choices 5<n<12,5<n<12(with underline under the pointys),7<n<17,7<n<12

OpenStudy (anonymous):

HI :DD

OpenStudy (anonymous):

hi

OpenStudy (anonymous):

can you help?

OpenStudy (anonymous):

larryboxaplenty

OpenStudy (anonymous):

someone please

OpenStudy (anonymous):

help please

OpenStudy (ranga):

The sum of two sides of a triangle is always GREATER than the third side. Therefore, 5 + 12 > n 17 > n or n < 17

OpenStudy (anonymous):

would that be a then?

OpenStudy (anonymous):

5<n<12?

OpenStudy (wolf1728):

According to this site http://www.1728.org/trianinq.htm Third side must be greater than 7 and less than 17 7 < third side < 17

OpenStudy (superdavesuper):

@ranga but n may NOT be the longest side, it could be shorter than 12 too. So it should be 7<n<17

OpenStudy (anonymous):

gotcha, thanks

OpenStudy (ranga):

Yes, true.

OpenStudy (wolf1728):

Agreed superdave

OpenStudy (superdavesuper):

Welcome n thank you all for the medal :)

OpenStudy (ranga):

I saw "write AN inequality" and so I just gave one inequality. Did not look at the answer choices posted later.